OpenStudy (anonymous):

Identify the vertex, focus, and directrix of the parabola with the equation x^2 – 6x – 8y + 49 = 0

OpenStudy (anonymous):

vertex (3,5)

OpenStudy (anonymous):

we have formula to figure out the focus and directrix. temporarily cannot remember. let me check

OpenStudy (anonymous):

yes. focus (3,10)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

like can you show me step by step please because i have a bunch of other problems that are the same :)

OpenStudy (anonymous):

to get vertex, you isolate y, take y', let y'=0 to get x, y of the vertex

OpenStudy (anonymous):

to get focus, you have formula x^2 =4py, your x is (x-3) your y = y-5---> p =10

OpenStudy (anonymous):

the directrix is y =-p = -10

OpenStudy (anonymous):

I think so. check your answer book. long time , i may not be right.

OpenStudy (anonymous):

thanks a lot :)

OpenStudy (anonymous):

you 're welcome